How does dx2-dy2 orbital have two nodal planes? How can they be in 45 degrees with x and y axis?

Grade:11

2 Answers

Umakant biswal
7 years ago
@ subhendra 
nodal plane is a plane in which the probability of finding a electron is zero. The 2s orbital has a nodal shell, whereas the 2pz orbital or 2p orbitals have a nodal plane.
so, when u will look the representation of these orbitals in crystal field theory , then u can easily detect that 
here dx2-dy2 orbital.. 
and x+y = 0 and x-y = 0 are the nodal planes, beacuse there are no probability of finding the electron their . 
and they are always lie at 45 degree with each other , as its the angular dependence of crystal field theory and beyond that the repulsion will be more .. 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..

 
Abhisek
4 years ago
Node is that region where wave function of electron changes it's sign from  + to -
Means it has to be zero
The angular part of dx2-y2 is propertional to 
X2-y2/r2=0
=>x2-y2=0
=>x=y or x=-y
Which are equation of plane 45 degrees with xaxis
:)Hope it helps